# GeoWarp: An automatically differentiable and GPU-accelerated implicit MPM framework for geomechanics based on NVIDIA Warp

### v1

@@ 1 Introduction
The material point method (MPM) [sulsky1994particle] is a continuum particle method that combines Lagrangian and Eulerian descriptions. It tracks the state of materials using a set of particles in a Lagrangian framework, while solving the governing equations on a background grid following an Eulerian approach. This hybrid Lagrangian–Eulerian formulation allows MPM to simulate large-deformation kinematics of history-dependent materials without suffering from mesh distortion. This capability makes MPM particularly well suited for modeling geomaterials (e.g. soils, rocks, and snow) that undergo large deformations. As a result, it has been increasingly adopted in computational geomechanics to model the behavior of geomaterials and their interactions with various objects and processes (e.g. [zabala2011progressive, yerro2015material, kularathna2021semi, dunatunga2015continuum, gaume2018dynamic, li2021three, seyedan2021solid, dunatunga2022modelling, jiang2022hybrid, jiang2023erosion, liang2023revealing, zhao2023coupled, yu2024semi, zhao2024mapped]).
@@
Most existing MPM implementations employ explicit time integration schemes—particularly the explicit Euler method—due to their algorithmic simplicity, ease of implementation, and suitability for dynamic problems. Accordingly, explicit MPM has been widely adopted in open-source codes such as Uintah-MPM [germain2000uintah], NairnMPM [nairn2015], CB-Geo MPM [kumar2019scalable], Karamelo [de2021karamelo], fMPM-solver [wyser2020fast], GeoTaich [shi2024geotaichi], Anura3D [Anura3D2025], MaterialPointSolver.jl [huo2025high], and Matter [blatny2025matter]. Despite their advantages, explicit schemes are only conditionally stable and impose severe limitations on the time step size. Furthermore, explicit methods do not allow control over numerical error per time step. Due to these reasons, explicit MPM is undesirable for quasi-static or long-term problems prevalent in geomechanics [lacroix2020life].
@@
Implicit time integration schemes, in contrast, are unconditionally stable and permit error control at each time step. These features make them more suitable for quasi-static or long-term processes, and implicit solvers have been widely employed in geomechanics (e.g. [young2009liquefaction, borja2010continuum, choo2015stabilized, choo2016hydromechanical, choo2018cracking]). However, implementing implicit MPM is considerably more challenging, as it requires solving a nonlinear system of equations at every time step. This is typically achieved using Newton’s method, which in turn demands accurate derivation and implementation of the Jacobian matrix to ensure convergence. Deriving the Jacobian matrix is particularly tedious and error-prone when the formulation involves complex constitutive models, as the consistent tangent operator—defined as the derivative of stress with respect to strain consistent with the time integration algorithm—should also be computed [simo1985consistent]. In the context of MPM, only a limited number of studies have pursued implicit formulations (e.g. [guilkey2003implicit, beuth2011solution, wang2016development, charlton2017igimp, coombs2020ample, coombs2020lagrangian, zhao2020stabilized, chandra2024mixed]). Among these, AMPLE [coombs2020ample] provides an open-source implementation of implicit MPM with clear documentation and example applications. However, as noted by the authors, AMPLE is primarily intended for conceptual demonstration and is not designed for high-performance computing or large-scale simulation. To date, no open-source, high-performance implementation of implicit MPM tailored specifically for computational geomechanics has been made available.
@@
Automatic differentiation (AD) offers a promising pathway toward implementing implicit solvers and enabling differentiable simulation. By constructing a computational graph during the forward pass and applying the chain rule in reverse during the backward pass, AD allows for the exact and automatic computation of gradients. This capability has significantly impacted simulation-based fields, including computational mechanics, where AD is increasingly used to simplify the implementation of nonlinear solvers. In the context of the finite element method (FEM), recent studies have demonstrated that AD can greatly reduce development effort by eliminating the need to manually derive and implement Jacobian matrices for complex, path-dependent constitutive models (e.g. [vigliotti2021automatic, xue2023jax, dummer2024robust]). Beyond implementation efficiency, AD also enables gradient-based optimization workflows, including inverse analysis, control, and learning. In the MPM community, several recent frameworks have successfully integrated AD with explicit MPM solvers. For example, ChainQueen [hu2019chainqueen], DiffTaichi [hu2019difftaichi], and PlasticineLab [huang2021plasticinelab] demonstrate differentiable simulations for soft robotics and control applications by embedding AD into GPU-accelerated MPM engines. However, these frameworks are primarily built for dynamic simulations using explicit time integration. The extension of AD to implicit MPM remains largely unexplored, despite its potential to simplify the implementation of Newton-type solvers and support inverse modeling in geomechanical problems.
@@
Leveraging recent advances in AD, here we introduce GeoWarp—an open-source, high-performance framework for implicit MPM tailored to computational geomechanics built on NVIDIA Warp [warp2022]. Warp is a Python-based simulation platform that supports AD and just-in-time compilation for efficient execution on both CPUs and GPUs. Its differentiable kernel-based programming model enables automatic and accurate gradient computation, while its Python interface facilitates rapid development. These features make Warp a suitable foundation for constructing differentiable solvers for large-scale mechanics problems.
@@
To address the computational cost typically associated with AD, we introduce a Jacobian construction strategy that leverages the sparsity pattern inherent in MPM. By exploiting the locality of particle–grid interactions, the method reduces reverse-mode evaluations to a small, fixed number of backward passes independent of problem size. This approach enables efficient large-scale simulation on modern GPUs, even for three-dimensional problems. Beyond forward simulation, the automatic differentiability of GeoWarp facilitates inverse modeling tasks such as material parameter identification and integration with gradient-based optimization and learning-based methods.
@@ 2 Material point method formulation
This section outlines the implicit MPM formulation adopted in this work. We begin by stating the initial–boundary value problem governing large-deformation mechanics, followed by its spatial discretization using the material point method. The resulting nonlinear system is then solved using a fully implicit scheme with Newton’s method. For brevity, we focus on describing the essential components of MPM, referring readers to comprehensive references for further details [jiang2016material, zhang2016material, fern2019material, nguyen2023material].
@@ 2.1 Problem statement
where $\bm{u}_{0}$ denotes the initial displacement field. To close the formulation, a constitutive law relating the Cauchy stress tensor $\bm{\sigma}$ to the deformation gradient $\bm{F}$ should be specified. In this study, we consider several constitutive models commonly used in geomechanics. For brevity, we omit their detailed formulations and refer the reader to standard references on constitutive models (e.g. [de2011computational, borja2013plasticity]).
@@ 2.2 Implicit MPM discretization
For the MPM weighting functions, the original MPM formulation [sulsky1994particle] employs linear finite element shape functions; however, such functions are prone to interpolation errors when particles cross cell boundaries. To alleviate this issue, several enhanced interpolation schemes have been developed, including the Generalized Interpolation Material Point (GIMP) [bardenhagen2004generalized] method, Convected Particle Domain Interpolation (CPDI) [sadeghirad2011convected, sadeghirad2013second], B-spline-based approaches [steffen2008examination, steffen2008analysis], and Moving Least Squares (MLS) interpolation [hu2018moving]. In this work, we adopt the contiguous GIMP (cpGIMP) scheme [wallstedt2008evaluation], which has been widely used in conjunction with implicit MPM formulations [charlton2017igimp, coombs2020ample, coombs2020lagrangian]. The one-dimensional weighting function and its gradient follow the formulation presented by Coombs and Augarde [coombs2020ample]. In multiple dimensions, the weighting functions and their gradients are constructed as tensor products of the corresponding one-dimensional components along each spatial direction.
@@
$-\int_{\Omega}\operatorname{\nabla^{s}}\bm{\eta}:\bm{\sigma}\,\mathrm{d}v+\int_{\Omega}\bm{\eta}\cdot\rho\bm{g}\,\mathrm{d}v+\int_{\partial\Omega_{\bm{t}}}\bm{\eta}\cdot\hat{\bm{t}}\,\mathrm{d}a=0,$ (8)
@@
The convergence behavior of Newton’s method is highly sensitive to the accuracy of the Jacobian matrix. In computational mechanics, the primary difficulty in evaluating the Jacobian lies in computing the consistent tangent operator, defined as the derivative of the incremental stress with respect to the incremental strain (or an equivalent measure) [simo1985consistent]. For relatively simple constitutive models, this operator can be derived analytically. However, for complex models commonly used in geomechanics, the derivation and implementation of consistent tangents is often tedious, error-prone, and model-specific. To overcome this limitation, we leverage AD to compute Jacobian matrices in a manner that is accurate, general, and fully automated. The details of our AD-based approach are presented in the following section.

### v2

@@ 1 Introduction
The material point method (MPM) [1] is a continuum particle method that combines Lagrangian and Eulerian descriptions. It tracks the state of materials using a set of particles in a Lagrangian framework, while solving the governing equations on a background grid following an Eulerian approach. This hybrid Lagrangian–Eulerian formulation allows MPM to simulate large-deformation kinematics of history-dependent materials without suffering from mesh distortion. This capability makes MPM particularly well suited for modeling geomaterials (e.g. soils, rocks, and snow) that undergo large deformations. As a result, it has been increasingly adopted in computational geomechanics to model the behavior of geomaterials and their interactions with various objects and processes (e.g. [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]).
@@
Most existing MPM implementations employ explicit time integration schemes—particularly the explicit Euler method—due to their algorithmic simplicity, ease of implementation, and suitability for dynamic problems. Accordingly, explicit MPM has been widely adopted in open-source codes such as Uintah-MPM [16], NairnMPM [17], CB-Geo MPM [18], Karamelo [19], fMPM-solver [20], GeoTaich [21], Anura3D [22], MaterialPointSolver.jl [23], and Matter [24]. Despite their advantages, explicit schemes are only conditionally stable and impose severe limitations on the time step size. Furthermore, explicit methods do not allow control over numerical error per time step. Due to these reasons, explicit MPM is undesirable for quasi-static or long-term problems prevalent in geomechanics [25].
@@
Implicit time integration schemes, in contrast, are unconditionally stable and permit error control at each time step. These features make them more suitable for quasi-static or long-term processes, and implicit solvers have been widely employed in geomechanics (e.g. [26, 27, 28, 29, 30]). However, implementing implicit MPM is considerably more challenging, as it requires solving a nonlinear system of equations at every time step. This is typically achieved using Newton’s method, which in turn demands accurate derivation and implementation of the Jacobian matrix to ensure convergence. Deriving the Jacobian matrix is particularly tedious and error-prone when the formulation involves complex constitutive models, as the consistent tangent operator—defined as the derivative of stress with respect to strain consistent with the time integration algorithm—should also be computed [31]. In the context of MPM, only a limited number of studies have pursued implicit formulations (e.g. [32, 33, 34, 35, 36, 37, 38, 39]). Among these, AMPLE [36] provides an open-source implementation of implicit MPM with clear documentation and example applications. However, as noted by the authors, AMPLE is primarily intended for conceptual demonstration and is not designed for high-performance computing or large-scale simulation. To date, no open-source, high-performance implementation of implicit MPM tailored specifically for computational geomechanics has been made available.
@@
Automatic differentiation (AD) offers a promising pathway toward implementing implicit solvers and enabling differentiable simulation. By constructing a computational graph during the forward pass and applying the chain rule in reverse during the backward pass, AD allows for the exact and automatic computation of gradients. This capability has significantly impacted simulation-based fields, including computational mechanics, where AD is increasingly used to simplify the implementation of nonlinear solvers. In the context of the finite element method (FEM), recent studies have demonstrated that AD can greatly reduce development effort by eliminating the need to manually derive and implement Jacobian matrices for complex, path-dependent constitutive models (e.g. [40, 41, 42]). Beyond implementation efficiency, AD also enables gradient-based optimization workflows, including inverse analysis, control, and learning. In the MPM community, several recent frameworks have successfully integrated AD with explicit MPM solvers. For example, ChainQueen [43], DiffTaichi [44], and PlasticineLab [45] demonstrate differentiable simulations for soft robotics and control applications by embedding AD into GPU-accelerated MPM engines. However, these frameworks are primarily built for dynamic simulations using explicit time integration. The extension of AD to implicit MPM remains largely unexplored, despite its potential to simplify the implementation of Newton-type solvers and support inverse modeling in geomechanical problems.
@@
Building on recent advances in AD, here we introduce GeoWarp—an open-source, high-performance framework for implicit MPM tailored to computational geomechanics built on NVIDIA Warp [46]. Warp is a Python-based simulation platform that supports AD and just-in-time compilation for efficient execution on both CPUs and GPUs. Its differentiable kernel-based programming model enables automatic and accurate gradient computation, while its Python interface facilitates rapid development. These features make Warp a suitable foundation for constructing differentiable solvers for large-scale mechanics problems.
@@
To address the computational cost typically associated with AD, we introduce a Jacobian construction strategy that leverages the sparsity pattern inherent in MPM. By exploiting the locality of particle–grid interactions, the method reduces reverse-mode evaluations to a small, fixed number of backward passes independent of problem size. This algorithm enables efficient large-scale simulation on modern GPUs, even for three-dimensional problems. Beyond forward simulation, the automatic differentiability of GeoWarp facilitates inverse modeling tasks such as material parameter identification and integration with gradient-based optimization and learning-based methods.
@@ 2 Material point method formulation
This section outlines the implicit MPM formulation adopted in this work. We begin by stating the initial–boundary value problem governing large-deformation mechanics, followed by its spatial discretization using the material point method. The resulting nonlinear system is then solved using a fully implicit scheme with Newton’s method. For brevity, we focus on describing the essential components of MPM, referring readers to comprehensive references for further details [47, 48, 49, 50].
@@ 2.1 Problem statement
where $\bm{u}_{0}$ denotes the initial displacement field. To close the formulation, a constitutive law relating the Cauchy stress tensor $\bm{\sigma}$ to the deformation gradient $\bm{F}$ should be specified. In this study, we consider several constitutive models commonly used in geomechanics. For brevity, we omit their detailed formulations and refer the reader to standard references on constitutive models (e.g. [51, 52]).
@@ 2.2 Implicit MPM discretization
For the MPM weighting functions, the original MPM formulation [1] employs linear finite element shape functions; however, such functions are prone to interpolation errors when particles cross cell boundaries. To alleviate this issue, several enhanced interpolation schemes have been developed, including the Generalized Interpolation Material Point (GIMP) [53] method, Convected Particle Domain Interpolation (CPDI) [54, 55], B-spline-based algorithms [56, 57], and Moving Least Squares (MLS) interpolation [58]. In this work, we adopt the contiguous GIMP (cpGIMP) scheme [59], which has been widely used in conjunction with implicit MPM formulations [35, 36, 37]. The one-dimensional weighting function and its gradient follow the formulation presented by Coombs and Augarde [36]. In multiple dimensions, the weighting functions and their gradients are constructed as tensor products of the corresponding one-dimensional components along each spatial direction.
@@
$\int_{\Omega}\operatorname{\nabla^{s}}\bm{\eta}:\bm{\sigma}\,\mathrm{d}v+\int_{\Omega}\bm{\eta}\cdot\rho\bm{g}\,\mathrm{d}v+\int_{\partial\Omega_{\bm{t}}}\bm{\eta}\cdot\hat{\bm{t}}\,\mathrm{d}a=0,$ (8)
@@
The convergence behavior of Newton’s method is highly sensitive to the accuracy of the Jacobian matrix. In computational mechanics, the primary difficulty in evaluating the Jacobian lies in computing the consistent tangent operator, defined as the derivative of the incremental stress with respect to the incremental strain (or an equivalent measure) [31]. For relatively simple constitutive models, this operator can be derived analytically. However, for complex models commonly used in geomechanics, the derivation and implementation of consistent tangents is often tedious, error-prone, and model-specific. To overcome this limitation, we leverage AD to compute Jacobian matrices in a manner that is accurate, general, and fully automated. The details of our AD-based approach are presented in the following section.